How much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 35% antifreeze?
Answer by ptaylor(2198)

Let x=amount of water needed
Since we are not disturbing the amount of pure antifreeze by adding or subtracting antifreeze, the amount of pure antifreeze before the mixture takes place(1 gal) has to equal the amount of pure antifreeze after the mixture takes place (0.35(1+x)). So our equation to solve is:
1=0.35(1+x) get rid of parens
1=0.35+0.35x subtract 0.35 from each side
1-0.35=0.35-0.35+0.35x collect like terms
0.35x=0.65 divide each side by 0.35
x=1.9 gal---------------------amont of water needed
CK
1=0.35*(1+1.9)=0.35*2.9
1~~~~~1
Hope this helps---ptaylor